OpenStudy (anonymous):

Help please:3x+y=5 solve for y

OpenStudy (kewlgeek555):

Hm. They don't give you something to go for x?

OpenStudy (kewlgeek555):

Well, if they don't: y = 5 - 3x.

OpenStudy (mathstudent55):

Since you are asked to solve for y, you want y by itslef on the left side. That means you look at the left side and see what you need to get rid of from the left side to have y by itself. You don't want the 3x there. Since 3x is being added to y, you need to subtract 3x from the left side. The rule is that whatever you do to one side of an equation, you must do to the other side also. That means subtract 3x from both sides: 3x + y = 5 -3x -3x --------------------- y = -3x + 5 Now it is solved for y.